# The Blueprints of Intelligence A Functional–Topological Foundation for Perception and Representation

### v4

@@ 1 Introduction
$x_{0}$$r\mathcal{M}$PMBatteryECG
Figure 1: Different physical systems (PM, battery, ECG) generate signals that lie on compact, low-variability manifolds. This schematic illustrates the universal geometric structure shared by deterministic perceptual manifolds.
@@ 7.2 Batteries: Electrochemical Discharge Manifolds
Battery discharge curves from the NASA Ames dataset exhibit smooth, deterministic voltage trajectories governed by electrochemical kinetics and diffusion dynamics. We analyzed nn=XXX discharge cycles from battery BXXXX under constant current loads of XX A.
Despite progressive capacity fade due to aging, the discharge manifold remains compact at any given degradation state. The voltage profile $V(t)$ exhibits characteristic morphology: an initial plateau regime, a gradual nonlinear decay, and a sharp terminal cutoff—all arising from the thermodynamics of lithium intercalation and internal resistance growth.
@@ 7.2.1 Saturation of the Battery Manifold
Computing geometric metrics over subsets $X_{n}$ of increasing size, we observe rapid stabilization:
@@
This saturation occurs despite the non-stationary nature of battery aging. The key insight is that at any given cycle number, the admissible discharge curves form a compact set shaped by the current electrochemical state, even as this state evolves slowly over hundreds of cycles.
@@ 7.2.2 Physical Interpretation
The rapid saturation reflects bounded electrochemical variability: manufacturing tolerances in electrode composition, small variations in ambient temperature, and stochastic effects in lithium plating all contribute to finite, structured deviations from the nominal trajectory. The perceptual radius captures the envelope of all physically realizable discharge profiles under normal operating conditions.
@@ Appendix B.5 Compactness
Figure 11: Example RMS envelopes derived from synthetic AC traces (gray) compared with a representative real RMS signal (green). Synthetic RMS profiles are obtained by computing RMS from the AC waveform, using the exact same processing pipeline applied to the real electromechanical data. The green curve shows a random real AC RMS sample, included to illustrate the typical morphology observed in field recordings.

### v5

@@ 1 Introduction
[…]
Figure 1: Different physical systems (PM, battery, ECG) generate signals that lie on compact, low-variability manifolds. This schematic illustrates the universal geometric structure shared by such deterministic perceptual manifolds: each domain has its own manifold in $C^{0}([0,T])$, but all exhibit the same compact, bounded-variability geometry.
@@ 2 Related Work
2.1 Industrial Assets: Railway Point Machines
Railway point machines constitute a representative class of industrial electromechanical systems, characterized by high-dimensional operational signals and strong physical constraints. Recent work has demonstrated the effectiveness of deep learning–based approaches for scalable diagnosis and predictive maintenance of point machines using large-scale operational data Ci et al. (2025). However, these approaches primarily focus on classification and prognostics performance, rather than on the intrinsic geometric structure of the space of physically admissible signals, which is the focus of the present work.
2.2 Physiological Signals: Electrocardiograms
Electrocardiogram (ECG) signals have been extensively studied for arrhythmia detection, morphological analysis, and clinical decision support using both classical signal processing and machine learning techniques. Despite the maturity of this literature, most existing approaches emphasize classification accuracy, feature extraction, or diagnostic performance, rather than investigating the global geometric structure and saturation properties of the space of admissible ECG waveforms.
2.3 Electrochemical Systems: Battery Discharge Signals
Battery discharge and degradation trajectories have been widely analyzed in the context of state-of-health estimation, remaining useful life prediction, and aging analysis, often using empirical electrochemical models or data-driven methods. While these works successfully characterize temporal evolution and degradation trends, they typically do not address whether the set of physically admissible discharge profiles forms a compact geometric manifold with intrinsic saturation properties.
@@ 7.1.2 Saturation of the Simulated PM Manifold
7.1.3 Waveform Morphology and Simulation Coverage
Figure 6: Representative RMS instantaneous power envelopes for railway point machines. In green: randomly sampled real manoeuvres from an in-service machine, showing the idle–inrush–locking–idle organisation with a terminal impact transient. In grey: Monte Carlo waveforms generated by the physics-aware AC simulator, which reproduce the same qualitative morphology while exploring admissible parametric variability.
@@ 7.2 Batteries: Electrochemical Discharge Manifolds
Battery discharge curves from the NASA Ames Prognostics Center of Excellence (PCoE) lithium-ion battery dataset exhibit smooth, deterministic voltage trajectories governed by electrochemical kinetics, diffusion processes, and internal resistance effects. We analyzed a total of n=2,794n=2{,}794 individual discharge cycles collected from $34$ commercial lithium-ion cells (B0005–B0056), spanning early-life to significantly aged operating regimes.
Each discharge corresponds to a constant-current load, with currents ranging approximately between $2$ A and $4$ A depending on the experimental protocol. While the nominal capacity of each cell decreases gradually over hundreds of cycles due to aging, each individual discharge curve represents a physically constrained realization of the underlying electrochemical system at a given degradation state.
Despite long-term non-stationarity across cycles, the instantaneous discharge manifold at any fixed aging stage remains compact. The voltage profile $V(t)$ exhibits a characteristic morphology shared across cells: an initial transient, a quasi-linear plateau regime, a nonlinear decay phase, and a sharp terminal cutoff associated with lithium depletion and increased internal resistance.
@@ 7.2.1 Saturation of the Battery Manifold
Geometric metrics were computed over subsets $X_{n}$ of discharge curves with increasing sample size. We observe rapid saturation of all metrics:
@@
Figure 7: Saturation of the battery discharge manifold. The Hausdorff distance between $X_{n}$ and $X_{n/2}$, as well as global radius estimates, stabilize rapidly, indicating that additional discharge curves do not introduce new geometric structure beyond a small number of samples.
Importantly, this saturation is observed despite the non-stationary nature of battery aging across cycles. The result indicates that while aging induces a slow drift of the discharge manifold over time, the set of physically realizable voltage trajectories at any given cycle remains strongly constrained and compact.
@@ 7.2.2 Physical Interpretation
The observed geometric saturation reflects bounded electrochemical variability. Manufacturing tolerances in electrode composition, small variations in ambient temperature, measurement noise, and stochastic effects in lithium transport introduce structured but finite deviations from the nominal discharge trajectory.
From a geometric perspective, the perceptual radius captures the envelope of all physically admissible discharge profiles under constant-current operation. Aging acts primarily as a slow deformation of this manifold rather than an expansion of its intrinsic dimensionality, explaining the early saturation observed across cycles.